OpenStudy (anonymous):

a 2000kilograms truck is braked to a top in 15meters from an initial speed of 12meters per second . how much force was required ?

OpenStudy (anonymous):

Please help me to solve this problem in physics , i need it right now :(

OpenStudy (samgrace):

F = ma --> a = F/m v^2 = u^2 + 2as <--- kinematic equation Substitute a=f/m into kinematic equation v^2 = u^2 + 2s x F/m. Final velocity v equals 0 to fit the equation to the problem u = 12 (m/s) s = 15 (m) 0^2 = 144 + (2 x 15) x (F/2000) (-144 / 30) x 2000 = F therefore F = -96000. The negative means that the force opposite to truck's motion, i.e in the opposite x direction.